### v1

# Accurate estimation of feature importance faithfulness for tree models

Abstract
In this paper, we consider a perturbation-based metric of predictive faithfulness of feature rankings (or attributions) that we call PGI squared. When applied to decision tree-based regression models, the metric can be computed accurately and efficiently for arbitrary independent feature perturbation distributions. In particular, the computation does not involve Monte Carlo sampling that has been typically used for computing similar metrics and which is inherently prone to inaccuracies.
Moreover, we propose a method of ranking features by their importance for the tree model’s predictions based on PGI squared. Our experiments indicate that in some respects, the method may identify the globally important features better than the state-of-the-art SHAP explainer.
1 Introduction
One of the key challenges in deploying modern machine learning models in such areas as medical diagnosis lies in the ability to indicate why a certain prediction has been made. Such an indication may be of critical importance when a human decides whether the prediction can be relied on. This is one of the reasons various aspects of explainability of machine learning models have been the subject of extensive research lately (see, e.g., [3]).
For some basic types of models (e.g., single decision trees), the rationale behind a prediction is easy to understand by a human. However, predictions of more complex models (that offer much better accuracy, e.g., based on neural networks or decision tree ensembles) are also much more difficult to interpret. Accurate and concise explanations understandable to humans might not always exist. In such cases, it is still beneficial to have methods giving a flavor of what factors might have influenced the prediction the most.
Local feature attribution [^1] It should be contrasted with global feature attribution whose aim is to measure the feature’s total impact on the model. constitutes one of such general approaches. For a fixed input $x$ and the model’s prediction $f(x)$, each feature is assigned a weight “measuring” the feature’s impact on the prediction. Ideally, a larger (absolute) weight should correspond to a larger importance of the feature for the prediction.
Many local attribution methods have been proposed so far, e.g., [12, 14, 15]. They are often completely model-agnostic, that is, they can be defined for and used with any model which is only accessed in a black-box way for computing the attribution. In some of these cases, focusing on a particular model’s architecture can lead to much more efficient and accurate attribution computation algorithms (e.g., [10] for tree ensemble models).
With the large body of different attribution methods, it is not clear which one should one use, especially since there is not a single objective measure of their reliability. To help deal with this issue, several notions of explanation quality have been proposed in the literature, such as faithfulness (fidelity) [11, 17], stability [2], or fairness [4, 8]. [1] developed an open-source benchmark OpenXAI automatically computing variants of these measures for a number of proposed attribution methods.
Out of these notions, our focus in this paper is on faithfulness.
Roughly speaking, an attribution method deserves to be called faithful if the features deemed important by the attribution are truly important for the decision-making process of the model. The perturbation-based methods constitute one popular class of approaches to measuring faithfulness. Intuitively, perturbing features deemed impactful should generally lead to a significant change in the prediction. On the contrary, manipulating unimportant features should not make a big difference.
One concrete cleanly-defined perturbation-based faithfulness metric for regression problems is the prediction gap on important feature perturbation, PGI in short. PGI is a faithfulness measure of choice in the OpenXAI benchmark [1]. Having fixed some subset of important features $S$ derived from the obtained attribution, the prediction gap wrt. $S$ is defined as the expected absolute change in the prediction when an independent Gaussian noise is applied to each coordinate $x_{j}$, $j\in S$. PGI can be defined as either the prediction gap wrt. a fixed set of important features (e.g., $k$ top-scoring features for some fixed $k$ [8]), or an average prediction gap over many important feature sets (as in [1]). In particular, the PGI metric depends only on the ordering of features by importance induced by the attribution. Other perturbation-based methods include e.g., PGU [1], comprehensiveness, sufficiency [7], remove-and-retrain [9], deletion [13].
In practice, PGI and other methods using random perturbations are computed using the Monte Carlo method which is inherently prone to inaccuracies, unless a very large number of samples is used.
It is interesting to ask whether PGI or similar random perturbation-based faithfulness matrices can be computed more accurately, say, with the approximation error caused merely by the floating-point arithmetic’s rounding errors.
1.1 Our contribution
1.1.1 Exact computation
First, we consider random perturbation-based measurement of faithfulness from the point of view of efficient and accurate computation. Of course, it is unrealistic to assume that such a quantity can be computed exactly (or even beyond the Monte Carlo method) given only black-box access to the model and the perturbation distribution. Hence, we focus on a concrete model architecture: tree ensemble models. Tree ensemble models remain a popular choice among practitioners since they are robust, easy to tune, and fast to train.
Unfortunately, computing the prediction gap, as defined in [1] (wrt. a fixed set $S$ of perturbed features, see Section 2 for a formal definition), does not look very tractable given the rich structure that tree ensemble models offer. This is because therein, the expectation is taken over the absolute value of $f(x)-f(x^{\prime})$, where $x^{\prime}$ is $x$ with important features perturbed. The absolute value is not very mathematically convenient to work with when combined with expectations. On the other hand, dropping the absolute value here would not make much sense: a small random perturbation (positive or negative) of a feature could in principle lead to a large positive or negative prediction difference. These, in turn, could cancel out in expectation, making the metric close to zero even though the feature is visibly of high importance.
To deal with the tractability problem of measuring PGI on tree models, we propose a similar metric, $\text{PGI}^{2}$, where instead of taking the absolute value, we compute the expectation of the square prediction gap $(f(x)-f(x^{\prime}))^{2}$. This eliminates the cancellation problem outlined above. Of course, introducing the square can alter the evaluation of the feature’s importance wrt. PGI as defined previously. However, intuitively, it amplifies the big prediction differences, and at the same time marginalizes the small, which is a generally good property if we seek concise explanations.
We prove that for tree ensemble models with $n$ nodes in total, the squared prediction gap can be computed in $O(n^{2})$ time exactly for any fixed choice of subset of perturbed important features, assuming the cumulative distribution function of feature-wise random perturbations can be evaluated in constant time. That is, the obtained running time bound does not require that the random perturbations are Gaussian or their distributions are equal across features.
The OpenXAI benchmark calculates PGI by taking the average expected prediction gap obtained when perturbing $1,2,\ldots,d$ most important features (where $d$ denotes the total number of features; see  [1, Appendix A]). In such a case, $d$ different expectations of absolute prediction gap are estimated, each of them via Monte Carlo sampling. In our case, $\text{PGI}^{2}$ is defined as the average expected squared prediction gap when perturbing $1,2,\ldots,d$ most important features. Using our prediction gap algorithm, $\text{PGI}^{2}$ is computed exactly (assuming infinite precision arithmetic; in reality, up to the precision error incurred by the usage of floating-point arithmetic) in $O(n^{2}d)$ time.
Experiments.
We accompany our theoretical development with an experimental comparison of the results obtained using the exact algorithm and Monte Carlo sampling.
The experiments confirm that Monte Carlo simulation gives noticeably less accurate results within the allocated time (iteration) budget comparable to the running time of our algorithm. Specifically, in our experiments, the normalized mean absolute error (NMAE) is typically between $0.02$ and $0.3$. The error measured this way generally decreases when increasing the size of the model which follows from the quadratic complexity of our algorithm. It is also decreasing in the magnitude of perturbations used: sampling is generally better at estimating larger quantities, whereas tiny perturbations lead to tiny prediction gaps.
1.1.2 $\text{PG}^{2}$-based greedy feature ordering.
We also investigate the possibility of using our exact $\text{PG}^{2}$ algorithm as a base of a feature importance ranking algorithm. We stress that finding a feature ordering optimizing the $\text{PGI}^{2}$ metric is a highly non-trivial task, as the metric depends on the entire ordering of the features. Checking all the features’ permutations is infeasible in most cases.
We consider constructing the feature ranking using a greedy $\text{PG}^{2}$ heuristic: the $i$-th most important feature $f$ is chosen so that it optimizes the prediction gap together with the already chosen $i-1$ most important features plus $f$.
We next compare the greedy $\text{PG}^{2}$ ranking with the ranking produced using the popular SHAP feature attribution method for tree ensembles [10]. First of all, our experiments confirm that the greedy $\text{PG}^{2}$ ranking yields, on average, better $\text{PGI}^{2}$ scores than SHAP.
While that such a property holds is non-obvious, it is perhaps not very surprising either. This is why we also compare our ranking method with SHAP in terms of average conciseness. Specifically, we consider three natural metrics of averaged feature importance across the test data set, each of them based solely on the feature ordering and clearly promoting the most important features. For each of these metrics, we compare the entropy of the obtained averaged importances. Intuitively, a smaller entropy indicates that only few features are identified as critically important to the model.
Our experiments show that the greedy $\text{PG}^{2}$ ranking used with perturbations of sufficiently large standard deviation can lead to global importance scores with smaller entropy compared to SHAP.
2 Preliminaries
Let us denote by $f:\mathbb{R}^{d}\to\mathbb{R}$ the output function of the considered regression model. We sometimes also use $f$ to refer to the model itself. The input $x\in\mathbb{R}^{d}$ to $f$ is called a feature vector, and we denote by $x_{i}$ the value of the $i$-th feature of $x$. Thus, we identify the set of features with $[d]$.
Tree models.
When talking about decision trees, we assume them to be binary and based on single-value splits. That is, each non-leaf node $v$ of a decision tree $\mathcal{T}$ has precisely two children $a_{v},b_{v}$. It is also assigned a feature $q_{v}\in[d]$ and a threshold value $t_{v}\in\mathbb{R}$. Each leaf node $l\in\mathcal{T}$ is in turn assigned a value $y_{v}\in\mathbb{R}$. We denote by $\mathcal{L}(\mathcal{T})$ the set of leaves of the tree $\mathcal{T}$.
The output $f_{\mathcal{T}}(x)$ of the tree $\mathcal{T}$ is computed by following a root-leaf path in $\mathcal{T}$: at a non-leaf node $v\in\mathcal{T}$, we descend either to the child $a_{v}$ if $x_{q_{v}}<t_{v}$, or to $b_{v}$ otherwise. When a leaf $l$ is eventually reached, its value $y_{l}$ is returned. We denote by $n_{\mathcal{T}}$ the number of nodes in $\mathcal{T}$.
When considering tree ensemble models $(\mathcal{T})_{i=1}^{m}$, the output $f(x)$ of the model is simply the sum of outputs $f_{\mathcal{T}_{i}}(x)$ of its $m$ individual trees. We generally use $n=\sum_{i=1}^{m}n_{\mathcal{T}_{i}}$ to refer to the total size of the tree ensemble model.
Prediction gap on important feature perturbation.
Let a ranking $\pi\in[d]\to[d]$ be a permutation of features giving their importance from highest to lowest. Denote by $\pi[1..\ell]$ the set $\{\pi(1),\ldots,\pi(\ell)\}$, that is, the $\ell$ top ranking features in $\pi$. For any $x\in\mathbb{R}^{d}$ and $S\subseteq[d]$, let $\text{perturb}(x,S)$ be some distribution on feature vectors $x^{\prime}\in\mathbb{R}^{d}$ satisfying $x^{\prime}_{i}=x_{i}$ for all $i\notin S$. AKS+ [1] define:
$\mathrm{PGI}(x,\pi,k):=\mathbb{E}_{x^{\prime}\sim\text{perturb}(x,\pi[1..k])}\left[|f(x^{\prime})-f(x)|\right].$
Based on that, they define an averaged metric:
$\mathrm{PGI}(x,\pi):=\frac{1}{d}\sum_{k=1}^{d}\mathrm{PGI}(x,\pi,k).$
We analogously define the $\mathrm{PGI}^{2}$ metric:
$\mathrm{PGI}^{2}(x,\pi,k)$ $:=\mathbb{E}_{x^{\prime}\sim\text{perturb}(x,\pi[1..k])}\left[\left(f(x^{\prime})-f(x)\right)^{2}\right],$
$\mathrm{PGI}^{2}(x,\pi)$ $:=\frac{1}{d}\sum_{k=1}^{d}\mathrm{PGI}^{2}(x,\pi,k).$
3 Computing $\mathrm{PGI}^{2}$ for tree ensemble models
Consider distributions $\text{perturb}(x,S)$ such that ${x^{\prime}\sim\text{perturb}(x,S)}$ satisfies:
  1. 1.
    for every $i\in S$, $x^{\prime}_{i}=x_{i}+\delta_{i}$, where $\delta_{i}\sim\mathcal{D}_{i}$, and
  2. 2.
    all $\delta_{j}$ ($j\in S$) are independent random variables.
Moreover, we assume that for all $i\in S$, the c.d.f. $F_{\mathcal{D}_{i}}$ of $\mathcal{D}_{i}$ can be evaluated in $O(1)$ time.
In the following, we show how to compute
$\mathrm{PG}^{2}(x,S):=\mathbb{E}_{x^{\prime}\sim\text{perturb}(x,S)}\left[(f(x^{\prime})-f(x))^{2}\right]$ (1)
under the assumptions made in the case when $f$ is a tree ensemble model $(\mathcal{T})_{i=1}^{m}$. Note that the metric $\mathrm{PGI}^{2}(x,\pi)$ for a given ranking $\pi$ can be computed by running the algorithm  $d$ times for different subsets $S$ of the form $\pi[1..k]$.
Put $c=f(x)$. Let $x^{\prime}\sim\text{perturb}(x,S)$. Recall that ${f(x^{\prime})=\sum_{i=1}^{m}f_{\mathcal{T}_{i}}(x^{\prime})}$. For every leaf node $v$ of some $\mathcal{T}_{i}$, consider a random indicator variable $X_{v}$ such that $X_{v}=1$ iff $\mathcal{T}_{i}$ evaluates to the value $y_{v}$ of leaf $v$ given $x^{\prime}$ as input. In particular, for any $i=1,\ldots,m$, there is exactly one $v\in\mathcal{L}(\mathcal{T}_{i})$ such that $X_{v}=1$. Set $\mathcal{L}=\bigcup_{i=1}^{m}\mathcal{L}(\mathcal{T}_{i})$. Then:
$f(x^{\prime})=\sum_{v\in\mathcal{L}}X_{v}\cdot y_{v}.$
By linearity of expectation, we obtain:
$\mathrm{PG}^{2}(x,S)$ $=\mathbb{E}\left[\left(\sum_{v\in\mathcal{L}}X_{v}y_{v}-c\right)^{2}\right]$
$=c^{2}+\mathbb{E}\left[\left(\sum_{u\in\mathcal{L}}X_{u}y_{u}\right)\left(\sum_{v\in\mathcal{L}}X_{v}y_{v}-2c\right)\right]$
$=c^{2}+\sum_{\begin{subarray}{c}u,v\in\mathcal{L}\\ u\neq v\end{subarray}}y_{u}y_{v}\Pr[X_{u}=1\land X_{v}=1]$
$\qquad\qquad+\sum_{u\in\mathcal{L}}\Pr[X_{u}=1]\cdot y_{u}\cdot(y_{u}-2c).$
The above formula reduces computing $\mathrm{PG}^{2}(x,S)$ to calculating probabilities of the form $\Pr[X_{u}=1\land X_{v}=1]$ (where $u\neq v$) or of the form $\Pr[X_{u}=1]$. We focus on the former task as dealing with the latter is analogous but easier.
For a node $w$, let $Q_{w}$ be the set of features appearing on the root-$w$ path (in the tree whose $w$ is the node of). Observe that for $w\in\mathcal{L}$, $X_{w}=1$ holds iff for each feature $q\in Q_{w}$, $x^{\prime}_{q}$ falls into a certain interval $I_{w,q}$. Namely, if the root-to-$w$ path consists of nodes $w_{1},\ldots,w_{k}=w$, then $I_{w,q}$ can be constructed as follows. Start with the interval (,)($\infty$,$\infty$)( - ∞ , ∞ ), and follow the path downwards. For each encountered node $w_{i}$ splitting on $q$, intersect the current $I_{w,q}$ with either:
  • $(-\infty,t_{w_{i}})$ if $w_{i+1}=a_{w_{i}}$, or
  • $[t_{w_{i}},\infty)$ if $w_{i+1}=b_{w_{i}}$, or
  • $\emptyset$ otherwise.
It follows that $X_{u}=1\land X_{v}=1$ iff for all $q\in Q_{u}\cup Q_{v}$, $x^{\prime}_{q}\in I_{u,q}\cap I_{v,q}$. Hence, if we denote by $l_{u,v,q},r_{u,v,q}$ the respective endpoints of $I_{u,q}\cap I_{v,q}$, by the independence of all perturbations $\delta_{j}$, we obtain that $\Pr[X_{u}=1\land X_{v}=1]$ equals $\Pi(u,v)$, where
$\Pi(u,v)=\left(\prod_{q\in(Q_{u}\cup Q_{v})\setminus S}[x_{q}\in I_{u,q}\cap I_{v,q}]\right)\cdot\left(\prod_{q\in(Q_{u}\cup Q_{v})\cap S}\left(F_{\mathcal{D}_{q}}(r_{u,v,q}-x_{q})-F_{\mathcal{D}_{q}}(l_{u,v,q}-x_{q})\right)\right).$ (2)
Implemented naively, computing a single value $\Pi(u,v)$ takes $O(D_{u}+D_{v})$ time, where $D_{u},D_{v}$ are the depths of $u$ and $v$ in their respective trees. Indeed, note that by traversing the root-to-$u$ and root-to-$v$ paths, we can find all the relevant intervals $I_{u,q},I_{v,q}$ that the product $\Pi(u,v)$ depends on. Moreover, $\Pi(u,v)$ has at most $D_{u}+D_{v}$ factors, each of which can be evaluated in $O(1)$ time.
Computing the values $\Pi(u,v)$ through all $\Theta(n^{2})$ pairs $u,v\in\mathcal{L}$ could take $O(n^{2}D)$ time, where $D$ is a global bound on the depth of the trees in the ensemble.
However, the computation can be optimized to run in $O(n^{2})$ time by computing the probabilities in an adequate order. Namely, note that, say, the value $\Pi(a_{u},v)$ can be computed based on $\Pi(u,v)$ in $O(1)$ time. This is because the interval $I_{a_{u},q}$ can differ from the corresponding interval $I_{u,q}$ only for the feature $q=q_{u}$, and it can be obtained by splitting $I_{u,q}$ with the threshold $t_{u}$. Hence, only one factor of the product (2) has to be added or replaced in order to obtain $\Pi(a_{u},v)$ from $\Pi(u,v)$. An analogous trick applies to computing $\Pi(b_{u},v)$, $\Pi(u,a_{v})$, $\Pi(u,b_{v})$ from $\Pi(u,v)$. Consequently, by iterating through nodes $u$ in pre-order, and then, once ${u\in\mathcal{L}}$ is fixed, through nodes $v$ in pre-order (the order of processing trees does not matter), all the required values $\Pi(u,v)$ can be computed in $O(n^{2})$ total time.
Theorem 3.1.
Let $f$ be a tree ensemble model whose trees have $n$ nodes in total. Let $x\in\mathbb{R}^{d}$. Then for any $S\subseteq[d]$, $\mathrm{PG}^{2}(x,S)$ can be computed in $O(n^{2})$ time. In particular, given a feature ranking $\pi$, for any $k\in[d]$, $\mathrm{PGI}^{2}(x,\pi,k)$ can be computed in $O(n^{2})$ time, whereas $\mathrm{PGI}^{2}(x,\pi)$ can be computed in $O(n^{2}d)$ time.
4 Experiments
4.1 Datasets
Following the previous work (e.g., [1, 8]), in the experiments we assume that the (real) perturbations used for computing the prediction gap (either exactly, or approximately) come from the same normal distribution $\mathcal{D}=\mathcal{N}(0,\sigma^{2})$. This is why we need to assume that all the features have numerical values.
Using the same perturbation distribution for all the features requires the features be standardised by subtracting the mean and dividing by the standard deviation. This ensures that if a feature is perturbed by a noise variable drawn from $\mathcal{D}$, the perturbation will have a similar effect. More specifically, it prevents a perturbation from $\mathcal{D}$ being relatively significant for one feature and insignificant for another.
In our experiments we used two datasets:
  1. 1.
    Red Wine Quality Dataset [5] The dataset contains 11 features related to physico-chemical attributes of wine, all numerical and continuous. The task is to predict the score of a wine, which is an integer between 1 and 10. The task can be considered as both a regression and a classification task. The dataset contains 1 59915991\ 5991 599 examples.
  2. 2.
    California Housing [16] The dataset contains information from the 1990 Californian census on some characteristics of a given Californian county and the median house value in that county. There are 8 numerical characteristics, such as length or number of bedrooms, and one categorical characteristic - proximity to the ocean. For the reasons outlined before, we decided to drop this feature and use a modified dataset. The task is to predict the median value of the house. The dataset contains 20 6402064020\ 64020 640 examples.
4.2 Models
All tree ensembles are implemented in the XGBoost library [6]. For each of the two datasets, three models of the following types were trained:
  • Single tree – a single decision tree with max_depth set to 4.
  • Bigger – gradient boosted trees with max_depth set to 4 and number of trees limited to 40.
In each case, the data set was split 80:20 into training and test sets. In addition, for each model and its constraint, a grid search was performed over selected hyperparameters. The details of the training process can be found in Appendix B.
4.3 Comparing squared prediction gap algorithms
In this section we compare our algorithm computing the squared prediction gap (as defined in Equation (1)) for tree ensembles described in Section 3 and the Monte Carlo sampling-based method. Specifically, the latter algorithm, given $x$ and the set $S\subseteq[d]$ or perturbed features, evaluates $f(x^{\prime})$ for $i$ randomly sampled $x^{\prime}\sim\text{perturb}(x,S)$ (see Section 3) and records the average value of $(f(x^{\prime})-f(x))^{2}$. We will call $i$ the number of iterations in the following.
The subject of the comparison was the accuracy of the results produced by the sampling algorithm. Since the accuracy of the sampling algorithm increases with the number of iterations $i$, to make the comparison fair, the number $i$ was chosen so that the wall time of both algorithms was close. Both algorithms were implemented in Python and all the arithmetic operations in the algorithms were carried out using the numpy.float32 type.
The computations were carried out on a FormatServer THOR E221 (Supermicro) server equipped with two AMD EPYC 7702 64-Core processors and 512 GB of RAM with operation system Ubuntu 22.04.1 LTS.
The comparison between the algorithms was carried out in different settings. The variable parameters were:
  • Model type $m$ trained on a selected dataset. In our case, there were four models in total (Section 4.2) trained on two datasets (Section 4.1).
  • The standard deviation $\sigma$ of a Gaussian used to perturb a feature. In our experiment, the evaluated values were $\{0.01,0,03,0.1,0.3\}$.
[…]
$\text{NMAE}(y,\hat{y})=\frac{\sum_{j=1}^{N}\mid y_{j}-\hat{y_{j}}\mid}{\sum_{j=1}^{N}{\mid y_{j}\mid}}$
where $y_{j}$ is the correct value of the prediction gap calculated by the exact algorithm, $\hat{y_{j}}$ is the prediction gap calculated by the sampling algorithm. NMAE was used as opposed to the relative error to suppress the effect of relatively large floating-point errors present when operating on very small numbers.
The results obtained in the experiment – the error as a function of the number of iterations – are depicted using the “Mean” curve for $\sigma=0.3$ and Bigger model for the Red Wine Quality dataset in Figure 1 and for $\sigma=0.03$ and Bigger model for the Californian Housing dataset in Figure 2. Analogous figures for other settings $(m,\sigma)$ can be found in Appendix B.
Additionally, the figures contain separate error curves for some selected sizes $|S|$ of the perturbed features set. As can be observed for a given iteration, the NMAE error behaves similarly for all subset sizes. Overall, the difference between the exact and sampling algorithms decreases as the number of iterations in the sampling algorithm increases. This is indeed anticipated since Monte Carlo sampling yields better accuracy with iteration increase; in fact the relative error decreases proportionally to $1/\sqrt{i}$.
Figure 1: Results of precision comparison described in 4.3 for $\sigma=0.3$ and for Bigger model for Red Wine Quality dataset.
Figure 2: Results of precision comparison described in 4.3 for $\sigma=0.03$ and for Bigger model for California housing dataset.
For the fair comparison, we chose the number of iterations (out of $9$ considered as explained above) in the sampling algorithm that came closest to the running time of the exact algorithm for each setting. The chosen number of iterations and the associated NMAE are shown in Tables 1, 2, 3 4.
Table 1: Number of iterations needed to perform to match execution time of exact algorithm on single tree model for Wine Quality Dataset described in 4.3.
$\sigma$ Number of iterations Relative error
0.01 100 0.08
0.03 100 0.14
0.1 100 0.15
0.3 1000 0.05
Table 2: Number of iterations needed to perform to match execution time of exact algorithm on Bigger model for Wine Quality Dataset described in 4.3.
$\sigma$ Number of iterations Relative error
0.01 100 0.08
0.03 100 0.11
0.1 100 0.12
0.3 1000 0.04
Table 3: Number of iterations needed to perform to match execution time of exact algorithm on single tree model for Californian Housing dataset described in 4.3.
$\sigma$ Number of iterations Relative error
0.01 100 0.13
0.03 100 0.15
0.1 100 0.15
0.3 1000 0.04
Table 4: Number of iterations needed to perform to match execution time of exact algorithm on Bigger tree model for Californian Housing dataset described in 4.3.
$\sigma$ Number of iterations Relative error
0.01 500 0.3
0.03 500 0.14
0.1 1000 0.06
0.3 2000 0.03
We conclude that the “fair” comparison confirms that the Monte Carlo simulation gives noticeably less accurate results than our exact method. We observe that the error generally decreases when increasing the size of the model. This is justified by the fact that our algorithm’s running time increases quadratically with $n$, whereas a single iteration of the Monte Carlo method runs in time at most linear in $n$. As a result, the bigger the model size is, the more iterations the Monte Carlo method can perform within the same time budget. Moreover, the error decreases with the standard deviation of the distribution used to obtain perturbations. This is also expected: for smaller perturbations, the floating-point rounding errors are of higher significance.
Table 5: Number of iterations needed to perform to match execution time of exact algorithm on Bigger Californian Housing model described in 4.3.
$\sigma$ Number of iterations Relative error
0.01 500 0.30
0.03 500 0.14
0.1 1000 0.06
0.3 2000 0.03
5 Feature ranking using $\mathrm{PG}^{2}$
We propose a method for ranking the features of a model using the squared prediction gap called greedy $\text{PG}^{2}$. Let $x\in\mathbb{R}^{d}$. Suppose the feature perturbations $\delta_{i}$ all come from $\mathcal{D}=\mathcal{N}(0,\sigma)$. The ranking $\pi_{\mathrm{PG}}^{\sigma}(x)$ is constructed in the following inductive manner.
  1. 1.
    For every $i\in[d]$, compute $p_{i}=\mathrm{PG}^{2}(x,\{i\})$. Set the most important feature $\pi^{\sigma}_{\mathrm{PG}}(x)(1)$ to be the feature $i$ for which $p_{i}$ is the highest.
  2. 2.
    Assume $\pi^{\sigma}_{\mathrm{PG}}(x)(1),...,\pi^{\sigma}_{\mathrm{PG}}(x)(l)$ are already computed. Let ${S=\{\pi^{\sigma}_{\mathrm{PG}}(x)(1),...,\pi^{\sigma}_{\mathrm{PG}}(x)(l)\}}$. For every $i\in[d]\setminus S$, compute $p_{i}:=\mathrm{PG}^{2}(x,S\cup\{i\})$. Let the next important feature $\pi^{\sigma}_{\mathrm{PG}}(x)(l+1)$ be the feature $i$ for which $p_{i}$ is the highest.
In the following, we compare the feature rankings obtained using the above method with the rankings obtained using the popular SHAP attribution framework [10].
For a given $x\in\mathbb{R}^{d}$, SHAP produces an attribution vector $(\phi)_{i=1}^{d}$ whose elements sum up to $f(x)-\mathbb{E}[f(x)]$ but can be both positive and negative [12]. We define the ranking $\pi_{\text{SHAP}}(x)$ to be obtained by sorting the features $[d]$ by the absolute value of $\phi_{i}$, from highest to lowest. This is justified by the fact that the predictions $f(x)$ can be smaller than the expected value; in such a case, all the attributions $\phi_{i}$ might be negative.
5.1 Global comparison vs SHAP wrt. $\mathrm{PGI}^{2}$
We perform a global comparison of $\pi^{\sigma}_{\mathrm{PG}}$ and $\pi_{\text{SHAP}}$ wrt. the $\mathrm{PGI}^{2}$ metric, that is, we check which ranking method gives higher $\mathrm{PGI}^{2}$ scores on average. More formally, fix a dataset/model combination $m$ and a perturbation distribution $\mathcal{D}^{\prime}=\mathcal{N}(0,\sigma^{\prime})$, where potentially $\sigma^{\prime}\neq\sigma$, as we did in Section 4.3. Let $X\subseteq{\mathbb{R}}^{d}$ be the corresponding dataset used in $m$. Then, define the average $\mathrm{PGI}^{2}$ scores as
$\overline{{\mathrm{PGI}}^{2}}(\pi_{\text{SHAP}})$ $=\frac{1}{|X|}\sum_{x\in X}\mathrm{PGI}^{2}(x,\pi_{\text{SHAP}}(x)),$
$\overline{{\mathrm{PGI}}^{2}}(\pi_{\mathrm{PGI}}^{\sigma})$ $=\frac{1}{|X|}\sum_{x\in X}\mathrm{PGI}^{2}(x,\pi_{\mathrm{PGI}}^{\sigma}(x)).$
Figures 345 and 6 present the comparison of the average scores $\overline{{\mathrm{PGI}}^{2}}(\pi_{\text{SHAP}})$ and $\overline{{\mathrm{PGI}}^{2}}(\pi_{\mathrm{PGI}}^{\sigma})$ for $\sigma=0.3$ as a function of $\sigma^{\prime}$. We observe that even if the parameter $\sigma$ used when computing the ranking $\pi_{\mathrm{PG}}^{\sigma}$ differs from the parameter $\sigma^{\prime}$ used for computing the $\mathrm{PGI}^{2}$ metrics, $\overline{{\mathrm{PGI}}^{2}}(\pi_{\mathrm{PGI}}^{\sigma})$ is consistently better than $\overline{{\mathrm{PGI}}^{2}}(\pi_{\text{SHAP}})$ for the Bigger models. For Single tree models, this is the case only for sufficiently large $\sigma^{\prime}$, but in particular for $\sigma^{\prime}=\sigma$. The figures presenting the results for other choices of $\sigma$ can be found in Appendix B.2.
Figure 3: Single tree model for the Wine Quality dataset.
Figure 4: Bigger model for the Wine Quality dataset.
Figure 5: Single tree model for the Housing dataset.
Figure 6: Bigger model for the Housing dataset
We conclude that the greedy optimization of the $\mathrm{PGI}^{2}(x,\pi)$ metric performs reasonably well, even though its correctness could only be guaranteed if it optimized the $\mathrm{PGI}^{2}(x,\pi,1)$ metric (see Section 2) instead.
5.2 Conciseness comparison vs SHAP
We also compare the global conciseness of the feature rankings $\pi_{\text{SHAP}}$ and $\pi_{\mathrm{PG}}^{\sigma}$. Intuitively, we would like to declare a feature ranking method more concise if it tends to produce rankings that identify the most important features of the model more clearly. For example, the information that two or three features of the model are identified as much more important than the others seems to say more about the model’s behavior in general than saying that all the features are equally important on average.
Our conciseness evaluation is performed using four different methods of converting a feature ranking $\pi(x)$ to a score vector $r^{\pi}(x):[d]\to\mathbb{R}_{\geq 0}$: geom, top1, top2, top3:
  • geom sets $(r^{\pi}(x))_{i}=1/2^{k}$ if $\pi(k)=i$, that is, the scores increase geometrically with the importance.
[…]
Having fixed (as before) the model-dataset combination $m$, $\sigma$, and a scoring method, we compute the aggregate feature scores $R_{i}$ through all data points $x$ from the test dataset $X$:
$R_{i}=\sum_{x\in X}(r^{\pi}(x))_{i}.$
Finally, the conciseness $C_{\pi}$ is computed as the entropy of the normalized vector $R_{1},\ldots,R_{d}$:
$C_{\pi}=\sum_{i=1}^{d}\frac{R_{i}}{\sum_{j}R_{j}}\log\left(\frac{R_{i}}{\sum_{j}R_{j}}\right).$
Intuitively, higher entropy means more uncertainty in the results.
Tables 6,78, and 9 present the results of our comparison for $\sigma\in\{0.1,0.3\}$ for Wine Quality models, and additionally for $\sigma=1$ for the Housing models. We observe that for the “single tree” models, the entropy that SHAP ranking yields is slightly smaller, for most of the scoring methods. However, there is not much uncertainty in the produced rankings since the models base their predictions only on a few features. Indeed, in the case of “single tree” models for both Wine Quality and Californian Housing datasets, only three features were present in trees.
For the bigger models consisting of numerous trees, all the features are used. In this more complex case, where the top features are not identified immediately, the greedy $\mathrm{PG}$-based ranking yields more concise rankings for the geom, top2 and top3 metrics for the Wine Quality dataset. For the Housing dataset, however, the obtained entropies for $\pi_{\text{SHAP}}$ are also consistently smaller than for the $\pi_{\mathrm{PG}}^{\sigma}$ rankings for $\sigma\in\{0.1,0.3\}$. Only for $\sigma=1$, the greedy $\mathrm{PG}^{2}$ method looks generally more concise.
Table 6: Entropy comparison between SHAP rankings and PG rankings for Bigger model for Wine Quality Dataset.
score SHAP PG $\sigma=0.1$ PG $\sigma=0.3$ PG $\sigma=1.0$
geom 3.245 3.285 3.157 2.740
top1 2.643 3.059 2.792 1.730
top2 3.120 3.232 3.019 2.402
top3 3.305 3.300 3.189 2.846
Table 7: Entropy comparison between SHAP rankings and PG rankings for Bigger model for Californian Housing Dataset.
score SHAP PG $\sigma=0.1$ PG $\sigma=0.3$ PG $\sigma=1.0$
geom 2.570 2.774 2.624 2.383
top1 1.748 2.324 1.979 0.534
top2 2.394 2.651 2.377 2.176
top3 2.634 2.834 2.676 2.622
Table 8: Entropy comparison between SHAP rankings and PG rankings for Single tree model for Wine Quality Dataset.
score SHAP PG, $\sigma=0.1$ PG, $\sigma=0.3$ PG, $\sigma=1.0$
geom 2.644 2.463 2.328 2.316
top1 1.725 1.482 1.372 1.305
top2 2.112 1.988 1.654 1.585
top3 2.432 2.207 2.073 2.012
Table 9: Entropy comparison between SHAP rankings and PG rankings for Single tree model for Californian Housing Dataset.
score SHAP PG $\sigma=0.1$ PG $\sigma=0.3$ PG $\sigma=1.0$
geom 2.094 2.197 2.175 2.138
top1 1.077 0.884 0.722 0.031
top2 1.549 1.580 1.602 1.633
top3 1.585 1.998 2.020 2.066
Appendix A Models training
The grid search was performed by using class $GridSearchCV$ from library XGBoost [6] with numbers of folds in cross-validation set to $5$. Additionally to parameters described in 4.2, the grid search included also parameters with the following possible values:
  • $eta$ with possible values in $\{0.01,0.1,0.2,0.3,0.4,0.5,0.6,0.7,0.8,0.9\}$
  • $subsample$ with possible values in ${0.01,0.1,0.2,0.3,0.4,0.5,0.6,0.7,0.8,0.9}$
Appendix B Additional experiments results
B.1 Results from precision experiments
Figure 7: Single tree model for Wine Quality dataset
Figure 8: Bigger model for Wine Quality dataset
Figure 9: Single tree model for Californian Housing dataset
Figure 10: Bigger model for Californian Housing dataset
B.2 Results from PGI comparison experiment
Figure 11: Single tree model for the Wine Quality dataset.
Figure 12: Bigger model for the Wine Quality dataset.
Figure 13: Single tree model for Californian Housing dataset
Figure 14: Bigger model for Californian Housing dataset

### v3

@@
-Accurate estimation of feature importance faithfulness for tree models
+AAAI Press Formatting Instructions for Authors Using — A Guide
 Abstract
 In this paper, we consider a perturbation-based metric of predictive faithfulness of feature rankings (or attributions) that we call PGI squared.
-When applied to decision tree-based regression models, the metric can be computed accurately and efficiently for arbitrary independent feature perturbation distributions.
+When applied to decision tree-based regression models, the metric can be computed exactly and efficiently for arbitrary independent feature perturbation distributions.
 In particular, the computation does not involve Monte Carlo sampling that has been typically used for computing similar metrics and which is inherently prone to inaccuracies.
-Moreover, we propose a method of ranking features by their importance for the tree model’s predictions based on PGI squared.
-Our experiments indicate that in some respects, the method may identify the globally important features better than the state-of-the-art SHAP explainer.
-1 Introduction
+As a second contribution, we proposed a procedure for constructing feature ranking based on PGI squared.
+Our results indicate the proposed ranking method is comparable to the widely recognized SHAP explainer, offering a viable alternative for assessing feature importance in tree-based models.
+Codehttps://github.com/rapicki/prediction-gap
+Extended versionhttps://arxiv.org/abs/2404.03426
+Introduction
 One of the key challenges in deploying modern machine learning models in such areas as medical diagnosis lies in the ability to indicate why a certain prediction has been made.
 Such an indication may be of critical importance when a human decides whether the prediction can be relied on.
-This is one of the reasons various aspects of explainability of machine learning models have been the subject of extensive research lately (see, e.g., [3]).
-For some basic types of models (e.g., single decision trees), the rationale behind a prediction is easy to understand by a human.
+This is one of the reasons various aspects of explainability of machine learning models have been the subject of extensive research lately (see, e.g., (Burkart and Huber 2021)).
+For some basic types of models (e.g., single decision trees, the rationale behind a prediction is easy to understand by a human.
 However, predictions of more complex models (that offer much better accuracy, e.g., based on neural networks or decision tree ensembles) are also much more difficult to interpret.
@@
 Ideally, a larger (absolute) weight should correspond to a larger importance of the feature for the prediction.
-Many local attribution methods have been proposed so far, e.g., [12, 14, 15].
+Many local attribution methods have been proposed so far, e.g., (Lundberg and Lee 2017; Plumb, Molitor, and Talwalkar 2018; Ribeiro, Singh, and Guestrin 2016).
 They are often completely model-agnostic, that is, they can be defined for and used with any model which is only accessed in a black-box way for computing the attribution.
-In some of these cases, focusing on a particular model’s architecture can lead to much more efficient and accurate attribution computation algorithms (e.g., [10] for tree ensemble models).
+In some of these cases, focusing on a particular model’s architecture can lead to much more efficient and accurate attribution computation algorithms (e.g., (Lundberg et al.
+2020) for tree ensemble models).
 With the large body of different attribution methods, it is not clear which one should one use, especially since there is not a single objective measure of their reliability.
-To help deal with this issue, several notions of explanation quality have been proposed in the literature, such as faithfulness (fidelity) [11, 17], stability [2], or fairness [4, 8].
-[1] developed an open-source benchmark OpenXAI automatically computing variants of these measures for a number of proposed attribution methods.
+To help deal with this issue, several notions of explanation quality have been proposed in the literature, such as faithfulness (fidelity) (Liu et al.
+2021; Yeh et al.
+2019), stability (Alvarez-Melis and Jaakkola 2018), or fairness (Balagopalan et al.
+2022; Dai et al.
+2022).
+Agarwal et al.
+(2022) developed an open-source benchmark OpenXAI automatically computing variants of these measures for a number of proposed attribution methods.
 Out of these notions, our focus in this paper is on faithfulness.
@@
 One concrete cleanly-defined perturbation-based faithfulness metric for regression problems is the prediction gap on important feature perturbation, PGI in short.
-PGI is a faithfulness measure of choice in the OpenXAI benchmark [1].
-Having fixed some subset of important features $S$ derived from the obtained attribution, the prediction gap wrt.
-$S$ is defined as the expected absolute change in the prediction when an independent Gaussian noise is applied to each coordinate $x_{j}$, $j\in S$.
-PGI can be defined as either the prediction gap wrt.
-a fixed set of important features (e.g., $k$ top-scoring features for some fixed $k$ [8]), or an average prediction gap over many important feature sets (as in [1]).
+PGI is a faithfulness measure of choice in the OpenXAI benchmark (Agarwal et al.
+2022).
+Having fixed some subset of important features $S$ derived from the obtained attribution, the prediction gap (PG) of $x$ wrt.
+$S$ is defined as:
+$\mathrm{PG}(x,S):=\mathbb{E}_{x^{\prime}\sim\text{perturb}(x,S)}\left[|f(x^{\prime})-f(x)|\right].$ (1)
+In (Agarwal et al.
+2022), the concrete perturbation method $\text{perturb}(x,S)$ is adding an independent Gaussian noise from $\mathcal{N}(0,\sigma^{2})$ to each coordinate $x_{j}$, $j\in S$.
+$\mathrm{PGI}(x)$ can be defined as either the prediction gap wrt.
+a fixed set of important features (e.g., $k$ top-scoring features for some fixed $k$ (Dai et al.
+2022)), or an average prediction gap over many important feature sets (as in (Agarwal et al.
+2022)).
 In particular, the PGI metric depends only on the ordering of features by importance induced by the attribution.
-Other perturbation-based methods include e.g., PGU [1], comprehensiveness, sufficiency [7], remove-and-retrain [9], deletion [13].
-In practice, PGI and other methods using random perturbations are computed using the Monte Carlo method which is inherently prone to inaccuracies, unless a very large number of samples is used.
-It is interesting to ask whether PGI or similar random perturbation-based faithfulness matrices can be computed more accurately, say, with the approximation error caused merely by the floating-point arithmetic’s rounding errors.
-1.1 Our contribution
-1.1.1 Exact computation
+Other perturbation-based methods include e.g., PGU (Agarwal et al.
+2022), comprehensiveness, sufficiency (DeYoung et al.
+2020), remove-and-retrain (Hooker et al.
+2019), deletion (Petsiuk, Das, and Saenko 2018).
+In practice, PGI and other methods based on random perturbations are computed using the Monte Carlo method which is inherently prone to inaccuracies, unless a large number of samples is used.
+It is interesting to ask whether PGI or similar random perturbation-based faithfulness metrics can be computed exactly, that is, via a closed-form calculation, so that any approximation error is caused merely by the floating-point arithmetic’s rounding errors.
+Our Contribution
+Exact Computation
 First, we consider random perturbation-based measurement of faithfulness from the point of view of efficient and accurate computation.
@@
 Tree ensemble models remain a popular choice among practitioners since they are robust, easy to tune, and fast to train.
-Unfortunately, computing the prediction gap, as defined in [1] (wrt.
-a fixed set $S$ of perturbed features, see Section 2 for a formal definition), does not look very tractable given the rich structure that tree ensemble models offer.
+Unfortunately, computing $\mathrm{PG}(x,S)$ (wrt.
+a fixed set $S$ of perturbed features, as in (Agarwal et al.
+2022), see (1)) does not look very tractable even given the rich structure that tree ensemble models offer.
 This is because therein, the expectation is taken over the absolute value of $f(x)-f(x^{\prime})$, where $x^{\prime}$ is $x$ with important features perturbed.
-The absolute value is not very mathematically convenient to work with when combined with expectations.
+The absolute value is not very mathematically convenient to work with when taking expectations.
 On the other hand, dropping the absolute value here would not make much sense: a small random perturbation (positive or negative) of a feature could in principle lead to a large positive or negative prediction difference.
 These, in turn, could cancel out in expectation, making the metric close to zero even though the feature is visibly of high importance.
-To deal with the tractability problem of measuring PGI on tree models, we propose a similar metric, $\text{PGI}^{2}$, where instead of taking the absolute value, we compute the expectation of the square prediction gap $(f(x)-f(x^{\prime}))^{2}$.
+Theoretical contribution: To deal with the tractability problem of measuring PGI on tree models, we propose a similar metric, $\mathrm{PGI}^{2}$, obtained from a slightly different squared prediction gap:
+$\mathrm{PG}^{2}(x,S):=\mathbb{E}_{x^{\prime}\sim\text{perturb}(x,S)}\left[(f(x^{\prime})-f(x))^{2}\right].$ (2)
 This eliminates the cancellation problem outlined above.
@@
 However, intuitively, it amplifies the big prediction differences, and at the same time marginalizes the small, which is a generally good property if we seek concise explanations.
-We prove that for tree ensemble models with $n$ nodes in total, the squared prediction gap can be computed in $O(n^{2})$ time exactly for any fixed choice of subset of perturbed important features, assuming the cumulative distribution function of feature-wise random perturbations can be evaluated in constant time.
+We prove that for tree ensemble models with $n$ nodes in total, the squared prediction gap $\mathrm{PG}^{2}(x,S)$ can be computed in $O(n^{2})$ time exactly for any fixed choice of a subset $S$ of perturbed important features, assuming the cumulative distribution function of feature-wise random perturbations can be evaluated in constant time.
 That is, the obtained running time bound does not require that the random perturbations are Gaussian or their distributions are equal across features.
-The OpenXAI benchmark calculates PGI by taking the average expected prediction gap obtained when perturbing $1,2,\ldots,d$ most important features (where $d$ denotes the total number of features; see  [1, Appendix A]).
-In such a case, $d$ different expectations of absolute prediction gap are estimated, each of them via Monte Carlo sampling.
-In our case, $\text{PGI}^{2}$ is defined as the average expected squared prediction gap when perturbing $1,2,\ldots,d$ most important features.
-Using our prediction gap algorithm, $\text{PGI}^{2}$ is computed exactly (assuming infinite precision arithmetic; in reality, up to the precision error incurred by the usage of floating-point arithmetic) in $O(n^{2}d)$ time.
-Experiments.
-We accompany our theoretical development with an experimental comparison of the results obtained using the exact algorithm and Monte Carlo sampling.
-The experiments confirm that Monte Carlo simulation gives noticeably less accurate results within the allocated time (iteration) budget comparable to the running time of our algorithm.
-Specifically, in our experiments, the normalized mean absolute error (NMAE) is typically between $0.02$ and $0.3$.
-The error measured this way generally decreases when increasing the size of the model which follows from the quadratic complexity of our algorithm.
-It is also decreasing in the magnitude of perturbations used: sampling is generally better at estimating larger quantities, whereas tiny perturbations lead to tiny prediction gaps.
-1.1.2 $\text{PG}^{2}$-based greedy feature ordering.
+We note that whereas our $\mathrm{PG}^{2}$ algorithm is quadratic, from the theoretical point of view, the important thing is that on tree models, a prediction gap-style quantity computation can be carried out exactly in polynomial time after all.
+Given a permutation $\pi$ ranking the $d$ features from the most to least important, the OpenXAI benchmark calculates $\mathrm{PGI}$ by taking the average prediction gap obtained when perturbing $k=1,2,\ldots,d$ most important features $\pi[1..k]=\{\pi(1),\ldots,\pi(k)\}$ (see (Agarwal et al.
+2022, Appendix A)).
+In such a case, $d$ different expectations of absolute prediction gap are estimated, each via Monte Carlo sampling.
+In our case, $\mathrm{PGI}^{2}(x)$ is analogously defined as:
+$\mathrm{PGI}^{2}(x,\pi):=\frac{1}{d}\sum_{k=1}^{d}\mathrm{PG}^{2}(x,\pi[1..k]).$ (3)
+Using our prediction gap algorithm, $\mathrm{PGI}^{2}$ is computed exactly (assuming infinite precision arithmetic; in reality, up to the precision error incurred by the usage of floating-point arithmetic) in $O(n^{2}d)$ time.
+Experiments: In our experiments, we evaluated different methods for calculating the Prediction Gap.
+Our study encompassed the exact algorithm as well as two sampling techniques integration techniques: Monte Carlo (MC) and Quasi-Monte Carlo (QMC).
+With the increase in iteration count, the outputs of MC and QMC visibly converged to the output of our exact algorithm; this confirms good numerical stability of our approach.
+When allocating an equivalent time budget across all methods, our findings revealed an advantage over MC and QMC, with QMC demonstrating a slight edge.
+Specifically, the Normalized Mean Absolute Error (NMAE) for MC was 0.13 for single models and decreased to 0.01 for bigger models.
+In comparison, QMC exhibited NMAEs of approximately 0.05 for single models and 0.002 for bigger ones.
+These results underscore the efficacy of sampling methods in the context of computing the PG, particularly for more sophisticated model structures.
+$\text{PG}^{2}$-Based Greedy Feature Ordering.
 We also investigate the possibility of using our exact $\text{PG}^{2}$ algorithm as a base of a feature importance ranking algorithm.
@@
 Checking all the features’ permutations is infeasible in most cases.
-We consider constructing the feature ranking using a greedy $\text{PG}^{2}$ heuristic: the $i$-th most important feature $f$ is chosen so that it optimizes the prediction gap together with the already chosen $i-1$ most important features plus $f$.
-We next compare the greedy $\text{PG}^{2}$ ranking with the ranking produced using the popular SHAP feature attribution method for tree ensembles [10].
-First of all, our experiments confirm that the greedy $\text{PG}^{2}$ ranking yields, on average, better $\text{PGI}^{2}$ scores than SHAP.
+We consider constructing the feature ranking using a greedy $\mathrm{PG}^{2}$ heuristic: the $i$-th most important feature $\lambda$ is chosen so that it optimizes the squared prediction gap together with the already chosen $i-1$ most important features plus $\lambda$.
+We next compare the greedy $\mathrm{PG}^{2}$ ranking with the ranking produced using the popular SHAP feature attribution method for tree ensembles (Lundberg et al.
+2020).
+We observe that the most important features identified by the two methods generally differ, and the deviation is clearer if the applied perturbations are smaller.
+As far as faithfulness is considered, our experiments confirm that the greedy $\mathrm{PG}^{2}$ ranking yields, on average, better $\mathrm{PGI}^{2}$ scores than SHAP.
 While that such a property holds is non-obvious, it is perhaps not very surprising either.
-This is why we also compare our ranking method with SHAP in terms of average conciseness.
-Specifically, we consider three natural metrics of averaged feature importance across the test data set, each of them based solely on the feature ordering and clearly promoting the most important features.
-For each of these metrics, we compare the entropy of the obtained averaged importances.
-Intuitively, a smaller entropy indicates that only few features are identified as critically important to the model.
-Our experiments show that the greedy $\text{PG}^{2}$ ranking used with perturbations of sufficiently large standard deviation can lead to global importance scores with smaller entropy compared to SHAP.
-2 Preliminaries
+This is why we also compare the $\mathrm{PG}^{2}$-based ranking method with SHAP in terms of two different faithfulness metrics: feature randomization and remove-and-retrain (Hooker et al.
+2019).
+In the former case and some of our datasets, the $\mathrm{PG}^{2}$-based rankings (used with Gaussian perturbations for small standard deviations) performed better than SHAP, whereas in the others, the results were comparable.
+Wrt.
+the latter metric, SHAP achieved better performance.
+Preliminaries
 Let us denote by $f:\mathbb{R}^{d}\to\mathbb{R}$ the output function of the considered regression model.
@@
 We generally use $n=\sum_{i=1}^{m}n_{\mathcal{T}_{i}}$ to refer to the total size of the tree ensemble model.
-Prediction gap on important feature perturbation.
-Let a ranking $\pi\in[d]\to[d]$ be a permutation of features giving their importance from highest to lowest.
-Denote by $\pi[1..\ell]$ the set $\{\pi(1),\ldots,\pi(\ell)\}$, that is, the $\ell$ top ranking features in $\pi$.
-For any $x\in\mathbb{R}^{d}$ and $S\subseteq[d]$, let $\text{perturb}(x,S)$ be some distribution on feature vectors $x^{\prime}\in\mathbb{R}^{d}$ satisfying $x^{\prime}_{i}=x_{i}$ for all $i\notin S$.
-AKS+ [1] define:
-$\mathrm{PGI}(x,\pi,k):=\mathbb{E}_{x^{\prime}\sim\text{perturb}(x,\pi[1..k])}\left[|f(x^{\prime})-f(x)|\right].$
-Based on that, they define an averaged metric:
-$\mathrm{PGI}(x,\pi):=\frac{1}{d}\sum_{k=1}^{d}\mathrm{PGI}(x,\pi,k).$
-We analogously define the $\mathrm{PGI}^{2}$ metric:
-$\mathrm{PGI}^{2}(x,\pi,k)$ $:=\mathbb{E}_{x^{\prime}\sim\text{perturb}(x,\pi[1..k])}\left[\left(f(x^{\prime})-f(x)\right)^{2}\right],$
-$\mathrm{PGI}^{2}(x,\pi)$ $:=\frac{1}{d}\sum_{k=1}^{d}\mathrm{PGI}^{2}(x,\pi,k).$
-3 Computing $\mathrm{PGI}^{2}$ for tree ensemble models
-Consider distributions $\text{perturb}(x,S)$ such that ${x^{\prime}\sim\text{perturb}(x,S)}$ satisfies:
+Computing $\mathrm{PG}^{2}$ for Tree Ensemble Models
+Let $S\subseteq[d]$ be a subset of important features and $x\in\mathbb{R}^{d}$. In this section, we show our polynomial exact algorithm computing $\mathrm{PG}^{2}(x,S)$, as defined in (2).
+We consider distributions $\text{perturb}(x,S)$ such that ${x^{\prime}\sim\text{perturb}(x,S)}$ satisfies:
   1. 1.
-    for every $i\in S$, $x^{\prime}_{i}=x_{i}+\delta_{i}$, where $\delta_{i}\sim\mathcal{D}_{i}$, and
+    for every perturbed feature $i\in S$, $x^{\prime}_{i}=x_{i}+\delta_{i}$, where $\delta_{i}\sim\mathcal{D}_{i}$, and distribution $\mathcal{D}_{i}$ is fixed, and the c.d.f.
+$F_{\mathcal{D}_{i}}$ of $\mathcal{D}_{i}$ can be evaluated in $O(1)$ time.
   2. 2.
+    for every non-perturbed feature $i\notin S$, $x^{\prime}_{i}=x_{i}$,
+  3. 3.
     all $\delta_{j}$ ($j\in S$) are independent random variables.
-Moreover, we assume that for all $i\in S$, the c.d.f. $F_{\mathcal{D}_{i}}$ of $\mathcal{D}_{i}$ can be evaluated in $O(1)$ time.
-In the following, we show how to compute
-$\mathrm{PG}^{2}(x,S):=\mathbb{E}_{x^{\prime}\sim\text{perturb}(x,S)}\left[(f(x^{\prime})-f(x))^{2}\right]$ (1)
+The main goal of this section is to show how to compute
+$\mathrm{PG}^{2}(x,S):=\mathbb{E}_{x^{\prime}\sim\text{perturb}(x,S)}\left[(f(x^{\prime})-f(x))^{2}\right]$
 under the assumptions made in the case when $f$ is a tree ensemble model $(\mathcal{T})_{i=1}^{m}$.
-Note that the metric $\mathrm{PGI}^{2}(x,\pi)$ for a given ranking $\pi$ can be computed by running the algorithm  $d$ times for different subsets $S$ of the form $\pi[1..k]$.
-Put $c=f(x)$.
+Recall from Equation (3) that the metric $\mathrm{PGI}^{2}(x,\pi)$ for a given ranking $\pi$ can be computed by running the algorithm $d$ times for $d$ different subsets $S$ of the form $\pi[1..k]$.
+Put $c:=f(x)$.
 Let $x^{\prime}\sim\text{perturb}(x,S)$.
@@
   • $\emptyset$ otherwise.
-It follows that $X_{u}=1\land X_{v}=1$ iff for all $q\in Q_{u}\cup Q_{v}$, $x^{\prime}_{q}\in I_{u,q}\cap I_{v,q}$.
+It follows that the event $X_{u}=1\land X_{v}=1$ occurs iff for all $q\in Q_{u}\cup Q_{v}$, $x^{\prime}_{q}\in I_{u,q}\cap I_{v,q}$.
 Hence, if we denote by $l_{u,v,q},r_{u,v,q}$ the respective endpoints of $I_{u,q}\cap I_{v,q}$, by the independence of all perturbations $\delta_{j}$, we obtain that $\Pr[X_{u}=1\land X_{v}=1]$ equals $\Pi(u,v)$, where
-$\Pi(u,v)=\left(\prod_{q\in(Q_{u}\cup Q_{v})\setminus S}[x_{q}\in I_{u,q}\cap I_{v,q}]\right)\cdot\left(\prod_{q\in(Q_{u}\cup Q_{v})\cap S}\left(F_{\mathcal{D}_{q}}(r_{u,v,q}-x_{q})-F_{\mathcal{D}_{q}}(l_{u,v,q}-x_{q})\right)\right).$ (2)
+$\begin{split}\Pi(u,v)=\left(\prod\limits_{q\in(Q_{u}\cup Q_{v})\setminus S}[x_{q}\in I_{u,q}\cap I_{v,q}]\right)\cdot\\ \prod\limits_{q\in(Q_{u}\cup Q_{v})\cap S}\left(F_{\mathcal{D}_{q}}(r_{u,v,q}-x_{q})-F_{\mathcal{D}_{q}}(l_{u,v,q}-x_{q})\right).\end{split}$ (4)
 Implemented naively, computing a single value $\Pi(u,v)$ takes $O(D_{u}+D_{v})$ time, where $D_{u},D_{v}$ are the depths of $u$ and $v$ in their respective trees.
@@
 Consequently, by iterating through nodes $u$ in pre-order, and then, once ${u\in\mathcal{L}}$ is fixed, through nodes $v$ in pre-order (the order of processing trees does not matter), all the required values $\Pi(u,v)$ can be computed in $O(n^{2})$ total time.
+This idea is the realized in Algorithm 1, where, while the two nested pre-order traversals over $\mathcal{T}$ proceed, the endpoints of intervals $I_{\cdot,\cdot}$ are maintained using the arrays $l,r$, and the factors of (4) are maintained using the array factor.
+Algorithm 1 Computing $\Pi(u,v)$ for all leaf pairs $u,v$ given model $\mathcal{T}$, a feature vector $x$, and important features $S\subseteq[d]$
+$\forall_{q\in[d]}\quad l[q]=\infty,r[q]=\infty,\text{factor}[q]=1$
+2:procedure Compute-Prob($v,\text{prod}=1,u=\textbf{null}$)
+3:     if $v$ is non-root $\triangleright$ update the maintained arrays
+4:         $p:=$ parent of $v$ in $\mathcal{T}$
+5:         $q:=q_{p}$
+6:         $(l^{\prime},r^{\prime},f^{\prime}):=(l[q],r[q],\text{factor}[q])$
+7:         if $v=a_{p}$ $\triangleright$ $v$ is the left child of its parent
+8:              $r[q]:=\min(r[q],t_{p})$
+9:         else
+10:              $l[q]:=\max(l[q],t_{p})$
+11:         end if
+12:         if $q\in S$ $\triangleright$ See Equation (4)
+13:              $\text{factor}[q]:=F_{\mathcal{D}_{q}}(r[q]-x_{q})-F_{\mathcal{D}_{q}}(l[q]-x_{q})$
+14:         else
+15:              $\text{factor}[q]:=[x_{q}\geq l[q]\land x_{q}\leq r[q]]$
+16:         end if
+17:         $\text{prod}:=(\text{prod}/f^{\prime})\cdot\text{factor}[q]$
+18:     end if
+19:     if $v$ is a leaf then
+20:         if $u=\textbf{null}$ then
+21:              for $T\in\mathcal{T}$ $\triangleright$ inner loop through trees
+22:                  Compute-Prob($\text{root}(T),\text{prod},v$)
+23:              end for
+24:         $\triangleright$ both leaves $u$ and $v$ are fixed
+25:              $\Pi(u,v):=\text{prod}$
+26:         end if
+27:     $\triangleright$ descend
+28:         Compute-Prob($a_{v},\text{prod},u$)
+29:         Compute-Prob($b_{v},\text{prod},u$)
+30:     end if
+31:     if $v$ is non-root $\triangleright$ revert changes to $l,r,\text{factor}$
+32:         $(l[q],r[q],\text{factor}[q]):=(l^{\prime},r^{\prime},f^{\prime})$
+33:     end if
+34:end procedure
+35:
+36:for $T\in\mathcal{T}$ $\triangleright$ outer loop through trees
+37:     Compute-Prob($\text{root}(T)$)
+38:end for
 Theorem 0.1.
-Let $f$ be a tree ensemble model whose trees have $n$ nodes in total.
-Let $x\in\mathbb{R}^{d}$.
-Then for any $S\subseteq[d]$, $\mathrm{PG}^{2}(x,S)$ can be computed in $O(n^{2})$ time.
-In particular, given a feature ranking $\pi$, for any $k\in[d]$, $\mathrm{PGI}^{2}(x,\pi,k)$ can be computed in $O(n^{2})$ time, whereas $\mathrm{PGI}^{2}(x,\pi)$ can be computed in $O(n^{2}d)$ time.
-4 Experiments
-4.1 Datasets
-Following the previous work (e.g., [1, 8]), in the experiments we assume that the (real) perturbations used for computing the prediction gap (either exactly, or approximately) come from the same normal distribution $\mathcal{D}=\mathcal{N}(0,\sigma^{2})$.
+Let $f$ be a tree ensemble model whose trees have $n$ nodes in total. Let $x\in\mathbb{R}^{d}$. Then for any $S\subseteq[d]$, $\mathrm{PG}^{2}(x,S)$ can be computed in $O(n^{2})$ time.
+Experiments: Exact vs Monte Carlo Sampling
+In this section, we compare our algorithm computing the squared prediction gap (as defined in Equation (2)) for tree ensembles as described earlier, the regular Monte Carlo sampling-based method (MC) and Quasi Monte Carlo sampling method (QMC) [^2] The details of quasi-random generators used in our Quasi Monte Carlo implementation can be found in the Appendix..
+Specifically, the Monte Carlo methods, given $x$ and the set $S\subseteq[d]$ or perturbed features, evaluates $f(x^{\prime})$ for $i$ randomly sampled $x^{\prime}\sim\text{perturb}(x,S)$ and records the average value of $(f(x^{\prime})-f(x))^{2}$ over the samples.
+We will call $i$ the number of iterations in the following.
+Recall that the accuracy of Monte Carlo methods is supposed to increase with the number of iterations $i$.
+The subject of the comparison was hence the accuracy of the results produced by the sampling algorithms as a function of the number of iterations with the goal of deciding which method (exact, MC, or QMC) is preferable and under what conditions.
+Specifically, having fixed a model $f$, we looked at the average difference between the values $\mathrm{PG}^{2}(x,S)$ computed using our closed-form algorithm and the two considered sampling methods.
+Using the output of our algorithm as the “ground truth” requires the algorithm to be numerically stable.
+So an indirect goal of the comparison was to establish whether our algorithm has good numerical properties.
+Experimental Setup
+Datasets
+Following the previous work (e.g., (Agarwal et al.
+2022; Dai et al.
+2022)), in each experiment we assume that the (real) perturbations used for computing the prediction gap (either exactly, or approximately) come from the same normal distribution $\mathcal{D}=\mathcal{N}(0,\sigma^{2})$.
 This is why we need to assume that all the features have numerical values.
-Using the same perturbation distribution for all the features requires the features be standardised by subtracting the mean and dividing by the standard deviation.
+Using the same perturbation distribution for all the features requires the features be standardized by subtracting the mean and dividing by the standard deviation.
 This ensures that if a feature is perturbed by a noise variable drawn from $\mathcal{D}$, the perturbation will have a similar effect.
 More specifically, it prevents a perturbation from $\mathcal{D}$ being relatively significant for one feature and insignificant for another.
-In our experiments we used two datasets:
+In our experiments, we used the following datasets:
   1. 1.
-    Red Wine Quality Dataset [5] The dataset contains 11 features related to physico-chemical attributes of wine, all numerical and continuous.
+    Red Wine Quality (Cortez et al.
+2009).
+The dataset contains 11 features wine, all numerical and continuous.
 The task is to predict the score of a wine, which is an integer between 1 and 10.
-The task can be considered as both a regression and a classification task.
-The dataset contains 1 59915991\ 5991 599 examples.
+We considered it as a regression task.
+The dataset contains 1 59915991\ 5991 599 examples and has 11 features.
   2. 2.
-    California Housing [16] The dataset contains information from the 1990 Californian census on some characteristics of a given Californian county and the median house value in that county.
-There are 8 numerical characteristics, such as length or number of bedrooms, and one categorical characteristic - proximity to the ocean.
+    California Housing (Torgo 2023) The dataset contains information from the 1990 Californian census.
+There are 8 numerical characteristics, and one categorical - proximity to the ocean.
 For the reasons outlined before, we decided to drop this feature and use a modified dataset.
 The task is to predict the median value of the house.
-The dataset contains 20 6402064020\ 64020 640 examples.
-4.2 Models
-All tree ensembles are implemented in the XGBoost library [6]. For each of the two datasets, three models of the following types were trained:
-  • Single tree – a single decision tree with max_depth set to 4.
-  • Bigger – gradient boosted trees with max_depth set to 4 and number of trees limited to 40.
+The dataset contains 20 6402064020\ 64020 640 examples and has 8 features.
+  3. 3.
+    Parkinson Telemonitoring Data (Tsanas and Little 2009) The dataset contains 5 87558755\ 8755 875 voice measurements from Parkinson’s disease patients, collected at home.
+It includes 17 numerical features, after dropping 3 categorical columns(ID, age) with a task to predict UPDRS motor and total scores.
+Models
+All tree ensembles are implemented in the XGBoost library (Chen and Guestrin 2016). For each of the datasets, models of the following two types were trained:
+  • Single tree – one decision tree with max_depth up to 4.
+  • Bigger – gradient boosted trees with max_depth up to 4 and number of trees limited to 40.
 In each case, the data set was split 80:20 into training and test sets.
 In addition, for each model and its constraint, a grid search was performed over selected hyperparameters.
-The details of the training process can be found in Appendix B.
-4.3 Comparing squared prediction gap algorithms
-In this section we compare our algorithm computing the squared prediction gap (as defined in Equation (1)) for tree ensembles described in Section 3 and the Monte Carlo sampling-based method.
-Specifically, the latter algorithm, given $x$ and the set $S\subseteq[d]$ or perturbed features, evaluates $f(x^{\prime})$ for $i$ randomly sampled $x^{\prime}\sim\text{perturb}(x,S)$ (see Section 3) and records the average value of $(f(x^{\prime})-f(x))^{2}$.
-We will call $i$ the number of iterations in the following.
-The subject of the comparison was the accuracy of the results produced by the sampling algorithm.
-Since the accuracy of the sampling algorithm increases with the number of iterations $i$, to make the comparison fair, the number $i$ was chosen so that the wall time of both algorithms was close.
-Both algorithms were implemented in Python and all the arithmetic operations in the algorithms were carried out using the numpy.float32 type.
+The details of the training process can be found in the Appendix.
+Measurements
+The comparison was carried out in different settings. For a selected dataset and sampling method (MC/QMC), the variable parameters were:
+  • Model type $m$. Recall that, in our case, there were two model types for a fixed dataset.
+  • The standard deviation $\sigma$ of a Gaussian used to perturb a feature. We used the values $\{0.1,0.3,1.0\}$.
+[…]
+Implementation
+All algorithms were implemented in Python with the key parts including inference on the tree ensemble being implemented as a C++ package and all the arithmetic operations in the algorithms were carried out using the numpy.float32 type or C++ float type.
 The computations were carried out on a FormatServer THOR E221 (Supermicro) server equipped with two AMD EPYC 7702 64-Core processors and 512 GB of RAM with operation system Ubuntu 22.04.1 LTS.
-The comparison between the algorithms was carried out in different settings. The variable parameters were:
-  • Model type $m$ trained on a selected dataset. In our case, there were four models in total (Section 4.2) trained on two datasets (Section 4.1).
-  • The standard deviation $\sigma$ of a Gaussian used to perturb a feature. In our experiment, the evaluated values were $\{0.01,0,03,0.1,0.3\}$.
-[…]
-$\text{NMAE}(y,\hat{y})=\frac{\sum_{j=1}^{N}\mid y_{j}-\hat{y_{j}}\mid}{\sum_{j=1}^{N}{\mid y_{j}\mid}}$
-where $y_{j}$ is the correct value of the prediction gap calculated by the exact algorithm, $\hat{y_{j}}$ is the prediction gap calculated by the sampling algorithm.
-NMAE was used as opposed to the relative error to suppress the effect of relatively large floating-point errors present when operating on very small numbers.
-The results obtained in the experiment – the error as a function of the number of iterations – are depicted using the “Mean” curve for $\sigma=0.3$ and Bigger model for the Red Wine Quality dataset in Figure 1 and for $\sigma=0.03$ and Bigger model for the Californian Housing dataset in Figure 2.
-Analogous figures for other settings $(m,\sigma)$ can be found in Appendix B.
-Additionally, the figures contain separate error curves for some selected sizes $|S|$ of the perturbed features set.
-As can be observed for a given iteration, the NMAE error behaves similarly for all subset sizes.
+Results
+Accuracy
+To estimate the accuracy of sampling-based methods as a function of the iteration count, we computed the Normalised Mean Absolute Error (NMAE), defined as:
+$\text{NMAE}(y,\hat{y})=\frac{\sum_{j=1}^{N}\mid y_{j}-\hat{y_{j}}\mid}{\sum_{j=1}^{N}{\mid y_{j}\mid}},$
+where $y_{j}$ is the respective value $\mathrm{PG}(x_{j},S_{j})$ computed by the exact algorithm, and $\hat{y_{j}}$ is the respective estimate obtained by the sampling algorithm.
+As opposed to the more popular MAPE error, NMAE was used because it is well defined in cases when true prediction is 00, which is often the case (for example when the perturbation is small).
+The results obtained in the experiment – the error as a function of the number of iterations – are depicted using the “Mean Monte Carlo” and “Mean Quasi Monte Carlo” curves for $\sigma=0.3$ and Bigger model for the Red Wine Quality dataset in Figure 1.
+Analogous figures for other datasets, models and $(m,\sigma)$ can be found in the Appendix.
+Figure 1: NMAE for $\sigma=0.3$ and for Bigger model for Red Wine Quality dataset.
 Overall, the difference between the exact and sampling algorithms decreases as the number of iterations in the sampling algorithm increases.
-This is indeed anticipated since Monte Carlo sampling yields better accuracy with iteration increase; in fact the relative error decreases proportionally to $1/\sqrt{i}$.
-Figure 1: Results of precision comparison described in 4.3 for $\sigma=0.3$ and for Bigger model for Red Wine Quality dataset.
-Figure 2: Results of precision comparison described in 4.3 for $\sigma=0.03$ and for Bigger model for California housing dataset.
-For the fair comparison, we chose the number of iterations (out of $9$ considered as explained above) in the sampling algorithm that came closest to the running time of the exact algorithm for each setting.
-The chosen number of iterations and the associated NMAE are shown in Tables 1, 2, 3 4.
-Table 1: Number of iterations needed to perform to match execution time of exact algorithm on single tree model for Wine Quality Dataset described in 4.3.
-$\sigma$ Number of iterations Relative error
-0.01 100 0.08
-0.03 100 0.14
-0.1 100 0.15
-0.3 1000 0.05
-Table 2: Number of iterations needed to perform to match execution time of exact algorithm on Bigger model for Wine Quality Dataset described in 4.3.
-$\sigma$ Number of iterations Relative error
-0.01 100 0.08
-0.03 100 0.11
-0.1 100 0.12
-0.3 1000 0.04
-Table 3: Number of iterations needed to perform to match execution time of exact algorithm on single tree model for Californian Housing dataset described in 4.3.
-$\sigma$ Number of iterations Relative error
-0.01 100 0.13
-0.03 100 0.15
-0.1 100 0.15
-0.3 1000 0.04
-Table 4: Number of iterations needed to perform to match execution time of exact algorithm on Bigger tree model for Californian Housing dataset described in 4.3.
-$\sigma$ Number of iterations Relative error
-0.01 500 0.3
-0.03 500 0.14
-0.1 1000 0.06
-0.3 2000 0.03
-We conclude that the “fair” comparison confirms that the Monte Carlo simulation gives noticeably less accurate results than our exact method.
+This is indeed anticipated since MC sampling yields better accuracy with iteration increase; in fact, the relative error decreases proportionally to $1/\sqrt{i}$.
+Additionally, the error for the QMC method is significantly lower, especially for the lower count of iterations.
+Moreover, the consistent $\text{NMAE}(y,\hat{y})$ error drop with the increase of the iteration count of the sampling methods indicates that our closed-form algorithm is numerically stable.
+Efficiency
+The running time of the sampling methods depends, clearly, linearly on the iterations count.
+As a result, the decision about which method to choose should depend on the acceptable accuracy.
+To this end, for each of the considered settings (dataset and parameters), we identified the number of iterations (out of $13$ considered thresholds as explained above) in the respective sampling algorithm that came closest to the running time of the exact algorithm for each setting.
+These corresponding numbers of iterations and the associated NMAEs are shown in Table 1.
+A larger acceptable NMAE than the corresponding threshold value suggests using the respective sampling-based method, and for smaller NMAEs than that, the exact algorithm is preferable.
+Monte Carlo Quasi Monte Carlo
+$\sigma$ Iterations NMAE Iterations NMAE
+Single tree Red Wine Quality model
+0.1 100 0.132 100 0.04
+0.3 100 0.136 100 0.049
+1.0 100 0.122 100 0.054
+Bigger Red Wine Quality model
+0.1 4000 0.019 4000 0.003
+0.3 8000 0.014 8000 0.002
+1.0 15000 0.009 15000 0.001
+Single California Housing model
+0.1 100 0.158 100 0.051
+0.3 100 0.14 100 0.043
+1.0 100 0.129 100 0.049
+Bigger California Housing model
+0.1 4000 0.018 4000 0.002
+0.3 10000 0.011 10000 0.001
+1.0 15000 0.01 15000 0.001
+Single Parkinson Telemonitoring model
+0.1 500 0.071 100 0.037
+0.3 500 0.064 500 0.011
+1.0 500 0.085 100 0.092
+Bigger Parkinson Telemonitoring model
+0.1 6000 0.015 4000 0.004
+0.3 15000 0.009 15000 0.002
+1.0 20000 0.008 20000 0.002
+Table 1: Iterations needed to match the execution time of our exact algorithm, and the corresponding NMAEs.
+We observe that the vanilla Monte Carlo simulation gives noticeably less accurate results than our exact method with a comparable allocated time budget.
 We observe that the error generally decreases when increasing the size of the model.
-This is justified by the fact that our algorithm’s running time increases quadratically with $n$, whereas a single iteration of the Monte Carlo method runs in time at most linear in $n$.
-As a result, the bigger the model size is, the more iterations the Monte Carlo method can perform within the same time budget.
-Moreover, the error decreases with the standard deviation of the distribution used to obtain perturbations.
-This is also expected: for smaller perturbations, the floating-point rounding errors are of higher significance.
-Table 5: Number of iterations needed to perform to match execution time of exact algorithm on Bigger Californian Housing model described in 4.3.
-$\sigma$ Number of iterations Relative error
-0.01 500 0.30
-0.03 500 0.14
-0.1 1000 0.06
-0.3 2000 0.03
-5 Feature ranking using $\mathrm{PG}^{2}$
-We propose a method for ranking the features of a model using the squared prediction gap called greedy $\text{PG}^{2}$.
+This is justified by the fact that our algorithm’s running time increases quadratically with $n$, whereas a single iteration of the Monte Carlo and Quasi Monte Carlo method runs in time at most linear in $n$.
+As a result, the bigger the model size is, the more iterations the Monte Carlo and Quasi Monte Carlo methods can perform within the same time budget.
+Moreover, for Monte Carlo the error decreases with the standard deviation of the distribution used to obtain perturbations.
+Feature Ranking Using $\mathrm{PG}^{2}$
+We also study and evaluate a method for ranking the features of a model using the squared prediction gap called greedy $\text{PG}^{2}$.
 Let $x\in\mathbb{R}^{d}$.
-Suppose the feature perturbations $\delta_{i}$ all come from $\mathcal{D}=\mathcal{N}(0,\sigma)$.
+Suppose the feature perturbations $\delta_{i}$ all come from $\mathcal{D}=\mathcal{N}(0,\sigma^{2})$.
 The ranking $\pi_{\mathrm{PG}}^{\sigma}(x)$ is constructed in the following inductive manner.
-  1. 1.
-    For every $i\in[d]$, compute $p_{i}=\mathrm{PG}^{2}(x,\{i\})$. Set the most important feature $\pi^{\sigma}_{\mathrm{PG}}(x)(1)$ to be the feature $i$ for which $p_{i}$ is the highest.
-  2. 2.
-    Assume $\pi^{\sigma}_{\mathrm{PG}}(x)(1),...,\pi^{\sigma}_{\mathrm{PG}}(x)(l)$ are already computed.
+For $l=0,1,\ldots,d-1$, assume we have already computed the $l$ top features $\pi^{\sigma}_{\mathrm{PG}}(x)(1),\ldots,\pi^{\sigma}_{\mathrm{PG}}(x)(l)$.
 Let ${S=\{\pi^{\sigma}_{\mathrm{PG}}(x)(1),...,\pi^{\sigma}_{\mathrm{PG}}(x)(l)\}}$.
 For every $i\in[d]\setminus S$, compute $p_{i}:=\mathrm{PG}^{2}(x,S\cup\{i\})$.
-Let the next important feature $\pi^{\sigma}_{\mathrm{PG}}(x)(l+1)$ be the feature $i$ for which $p_{i}$ is the highest.
-In the following, we compare the feature rankings obtained using the above method with the rankings obtained using the popular SHAP attribution framework [10].
-For a given $x\in\mathbb{R}^{d}$, SHAP produces an attribution vector $(\phi)_{i=1}^{d}$ whose elements sum up to $f(x)-\mathbb{E}[f(x)]$ but can be both positive and negative [12].
+We define the next important feature $\pi^{\sigma}_{\mathrm{PG}}(x)(l+1)$ to be the feature $i$ for which $p_{i}$ is the highest.
+In the following, we compare the feature rankings obtained using the above method with the rankings obtained using the popular SHAP attribution framework (Lundberg et al.
+2020) for tree ensembles, concretely using the tree_path_dependent feature perturbation method.
+For a given $x\in\mathbb{R}^{d}$, SHAP produces an attribution vector $(\phi)_{i=1}^{d}$ whose elements sum up to $f(x)-\mathbb{E}[f(x)]$ but can be both positive and negative (Lundberg and Lee 2017).
 We define the ranking $\pi_{\text{SHAP}}(x)$ to be obtained by sorting the features $[d]$ by the absolute value of $\phi_{i}$, from highest to lowest.
 This is justified by the fact that the predictions $f(x)$ can be smaller than the expected value; in such a case, all the attributions $\phi_{i}$ might be negative.
-5.1 Global comparison vs SHAP wrt. $\mathrm{PGI}^{2}$
+A natural question arises whether this ranking method produces rankings similar or dissimilar to SHAP rankings.
+Table 2 shows the ratio of datapoints for which the top $k$ most important features are the same for $\pi^{\sigma}_{\mathrm{PG}}$ and $\pi_{\text{SHAP}}$.
+As we see, this ratio usually decreases with $k$, increases with $\sigma$, and is lower for the bigger models than for the single tree models.
+However, even the rankings for single tree models pick the same most important feature as SHAP less than 76% of the time.
+In general, the rankings produced are quite different, so it is worthwhile to study these differences and perform a comparative analysis of the rankings.
+$\sigma$ $k=1$ $k=2$ $k=3$
+Single tree Red Wine Quality model
+0.1 0.328 0.259 0.031
+0.3 0.412 0.591 0.428
+1.0 0.419 0.613 0.506
+Bigger Red Wine Quality model
+0.1 0.131 0.041 0.022
+0.3 0.212 0.094 0.056
+1.0 0.391 0.181 0.103
+Single California Housing model
+0.1 0.501 0.263 0.031
+0.3 0.599 0.569 0.349
+1.0 0.712 0.471 0.504
+Bigger California Housing model
+0.1 0.194 0.08 0.052
+0.3 0.3 0.122 0.056
+1.0 0.569 0.243 0.08
+Single Parkinson Telemonitoring model
+0.1 0.402 0.143 0.066
+0.3 0.569 0.32 0.237
+1.0 0.759 0.414 0.22
+Bigger Parkinson Telemonitoring model
+0.1 0.137 0.028 0.007
+0.3 0.225 0.061 0.009
+1.0 0.368 0.114 0.031
+Table 2: Ratio of datapoints having the same $k$ most important features in $\pi_{\mathrm{PG}}^{\sigma}$ as in $\pi_{\text{SHAP}}$
+Comparison vs SHAP wrt. $\mathrm{PGI}^{2}$
 We perform a global comparison of $\pi^{\sigma}_{\mathrm{PG}}$ and $\pi_{\text{SHAP}}$ wrt.
 the $\mathrm{PGI}^{2}$ metric, that is, we check which ranking method gives higher $\mathrm{PGI}^{2}$ scores on average.
-More formally, fix a dataset/model combination $m$ and a perturbation distribution $\mathcal{D}^{\prime}=\mathcal{N}(0,\sigma^{\prime})$, where potentially $\sigma^{\prime}\neq\sigma$, as we did in Section 4.3.
+More formally, fix a dataset/model combination $m$ and a perturbation distribution $\mathcal{D}^{\prime}=\mathcal{N}(0,(\sigma^{\prime})^{2})$, where potentially $\sigma^{\prime}\neq\sigma$, for perturbations used in computing $\mathrm{PG}^{2}(x,\cdot)$ in (3).
 Let $X\subseteq{\mathbb{R}}^{d}$ be the corresponding dataset used in $m$.
-Then, define the average $\mathrm{PGI}^{2}$ scores as
-$\overline{{\mathrm{PGI}}^{2}}(\pi_{\text{SHAP}})$ $=\frac{1}{|X|}\sum_{x\in X}\mathrm{PGI}^{2}(x,\pi_{\text{SHAP}}(x)),$
-$\overline{{\mathrm{PGI}}^{2}}(\pi_{\mathrm{PGI}}^{\sigma})$ $=\frac{1}{|X|}\sum_{x\in X}\mathrm{PGI}^{2}(x,\pi_{\mathrm{PGI}}^{\sigma}(x)).$
-Figures 345 and 6 present the comparison of the average scores $\overline{{\mathrm{PGI}}^{2}}(\pi_{\text{SHAP}})$ and $\overline{{\mathrm{PGI}}^{2}}(\pi_{\mathrm{PGI}}^{\sigma})$ for $\sigma=0.3$ as a function of $\sigma^{\prime}$.
-We observe that even if the parameter $\sigma$ used when computing the ranking $\pi_{\mathrm{PG}}^{\sigma}$ differs from the parameter $\sigma^{\prime}$ used for computing the $\mathrm{PGI}^{2}$ metrics, $\overline{{\mathrm{PGI}}^{2}}(\pi_{\mathrm{PGI}}^{\sigma})$ is consistently better than $\overline{{\mathrm{PGI}}^{2}}(\pi_{\text{SHAP}})$ for the Bigger models.
-For Single tree models, this is the case only for sufficiently large $\sigma^{\prime}$, but in particular for $\sigma^{\prime}=\sigma$.
-The figures presenting the results for other choices of $\sigma$ can be found in Appendix B.2.
-Figure 3: Single tree model for the Wine Quality dataset.
-Figure 4: Bigger model for the Wine Quality dataset.
-Figure 5: Single tree model for the Housing dataset.
-Figure 6: Bigger model for the Housing dataset
-We conclude that the greedy optimization of the $\mathrm{PGI}^{2}(x,\pi)$ metric performs reasonably well, even though its correctness could only be guaranteed if it optimized the $\mathrm{PGI}^{2}(x,\pi,1)$ metric (see Section 2) instead.
-5.2 Conciseness comparison vs SHAP
-We also compare the global conciseness of the feature rankings $\pi_{\text{SHAP}}$ and $\pi_{\mathrm{PG}}^{\sigma}$.
-Intuitively, we would like to declare a feature ranking method more concise if it tends to produce rankings that identify the most important features of the model more clearly.
-For example, the information that two or three features of the model are identified as much more important than the others seems to say more about the model’s behavior in general than saying that all the features are equally important on average.
-Our conciseness evaluation is performed using four different methods of converting a feature ranking $\pi(x)$ to a score vector $r^{\pi}(x):[d]\to\mathbb{R}_{\geq 0}$: geom, top1, top2, top3:
-  • geom sets $(r^{\pi}(x))_{i}=1/2^{k}$ if $\pi(k)=i$, that is, the scores increase geometrically with the importance.
-[…]
-Having fixed (as before) the model-dataset combination $m$, $\sigma$, and a scoring method, we compute the aggregate feature scores $R_{i}$ through all data points $x$ from the test dataset $X$:
-$R_{i}=\sum_{x\in X}(r^{\pi}(x))_{i}.$
-Finally, the conciseness $C_{\pi}$ is computed as the entropy of the normalized vector $R_{1},\ldots,R_{d}$:
-$C_{\pi}=\sum_{i=1}^{d}\frac{R_{i}}{\sum_{j}R_{j}}\log\left(\frac{R_{i}}{\sum_{j}R_{j}}\right).$
-Intuitively, higher entropy means more uncertainty in the results.
-Tables 6,78, and 9 present the results of our comparison for $\sigma\in\{0.1,0.3\}$ for Wine Quality models, and additionally for $\sigma=1$ for the Housing models.
-We observe that for the “single tree” models, the entropy that SHAP ranking yields is slightly smaller, for most of the scoring methods.
-However, there is not much uncertainty in the produced rankings since the models base their predictions only on a few features.
-Indeed, in the case of “single tree” models for both Wine Quality and Californian Housing datasets, only three features were present in trees.
-For the bigger models consisting of numerous trees, all the features are used.
-In this more complex case, where the top features are not identified immediately, the greedy $\mathrm{PG}$-based ranking yields more concise rankings for the geom, top2 and top3 metrics for the Wine Quality dataset.
-For the Housing dataset, however, the obtained entropies for $\pi_{\text{SHAP}}$ are also consistently smaller than for the $\pi_{\mathrm{PG}}^{\sigma}$ rankings for $\sigma\in\{0.1,0.3\}$.
-Only for $\sigma=1$, the greedy $\mathrm{PG}^{2}$ method looks generally more concise.
-Table 6: Entropy comparison between SHAP rankings and PG rankings for Bigger model for Wine Quality Dataset.
-score SHAP PG $\sigma=0.1$ PG $\sigma=0.3$ PG $\sigma=1.0$
-geom 3.245 3.285 3.157 2.740
-top1 2.643 3.059 2.792 1.730
-top2 3.120 3.232 3.019 2.402
-top3 3.305 3.300 3.189 2.846
-Table 7: Entropy comparison between SHAP rankings and PG rankings for Bigger model for Californian Housing Dataset.
-score SHAP PG $\sigma=0.1$ PG $\sigma=0.3$ PG $\sigma=1.0$
-geom 2.570 2.774 2.624 2.383
-top1 1.748 2.324 1.979 0.534
-top2 2.394 2.651 2.377 2.176
-top3 2.634 2.834 2.676 2.622
-Table 8: Entropy comparison between SHAP rankings and PG rankings for Single tree model for Wine Quality Dataset.
-score SHAP PG, $\sigma=0.1$ PG, $\sigma=0.3$ PG, $\sigma=1.0$
-geom 2.644 2.463 2.328 2.316
-top1 1.725 1.482 1.372 1.305
-top2 2.112 1.988 1.654 1.585
-top3 2.432 2.207 2.073 2.012
-Table 9: Entropy comparison between SHAP rankings and PG rankings for Single tree model for Californian Housing Dataset.
-score SHAP PG $\sigma=0.1$ PG $\sigma=0.3$ PG $\sigma=1.0$
-geom 2.094 2.197 2.175 2.138
-top1 1.077 0.884 0.722 0.031
-top2 1.549 1.580 1.602 1.633
-top3 1.585 1.998 2.020 2.066
+Then, for $\pi\in\{\pi_{\text{SHAP}},\pi_{\mathrm{PGI}}^{\sigma}\}$, define the average $\mathrm{PGI}^{2}$ scores as
+$\overline{{\mathrm{PGI}}^{2}}(\pi)$ $=\frac{1}{|X|}\sum_{x\in X}\mathrm{PGI}^{2}(x,\pi(x)).$
+Table 3 presents the average scores $\overline{{\mathrm{PGI}}^{2}}(\pi_{\text{SHAP}})$ and $\overline{{\mathrm{PGI}}^{2}}(\pi_{\mathrm{PGI}}^{\sigma})$ for $\sigma\in\{0.1,0.3,1.0\}$ as a function of $\sigma^{\prime}$ for Bigger models (an analogous table for Single models can be found in the Appendix).
+We observe that even if the parameter $\sigma$ used when computing the ranking $\pi_{\mathrm{PG}}^{\sigma}$ differs from the parameter $\sigma^{\prime}$ used for computing the $\mathrm{PGI}^{2}$ metrics, $\overline{{\mathrm{PGI}}^{2}}(\pi_{\mathrm{PGI}}^{\sigma})$ is almost always better than $\overline{{\mathrm{PGI}}^{2}}(\pi_{\text{SHAP}})$ for the Bigger models, with a slight exception in the ${\sigma^{\prime}=1.0,\sigma=0.1}$ pair case, when $\sigma,\sigma^{\prime}$ differ a lot.
+$\sigma^{\prime}$ SHAP $\sigma$ = $0.1$ $\sigma$ = $0.3$ $\sigma$ = $1.0$
+Bigger PGI Red Wine Quality model
+0.1 0.013 0.024 0.022 0.018
+0.3 0.051 0.068 0.078 0.067
+1.0 0.221 0.192 0.266 0.302
+Bigger PGI California Housing model
+0.1 3.346e+08 4.205e+08 3.841e+08 3.065e+08
+0.3 1.209e+09 1.297e+09 1.374e+09 1.259e+09
+1.0 7.807e+09 6.963e+09 8.266e+09 8.728e+09
+Bigger PGI Parkinson Telemonitoring model
+0.1 2.743 4.425 4.015 3.199
+0.3 6.341 9.458 11.878 10.143
+1.0 16.298 19.56 30.188 34.013
+Table 3: Greedy $\mathrm{PG}^{2}$ vs SHAP rankings wrt. $\mathrm{PGI}^{2}$.
+We conclude that the greedy optimization of the $\mathrm{PGI}^{2}(x,\pi)$ metric performs reasonably well, even though its correctness could only be guaranteed if it optimized the $\mathrm{PGI}(x,\{\pi_{1}\})$ metric (where only the single most important feature is perturbed) instead of (3).
+Comparison vs SHAP with Feature Removing
+We compare the rankings $\pi_{\text{SHAP}}$ and $\pi_{\mathrm{PG}}^{\sigma}$ using feature removing.
+Intuitively, removing information about the value of a highly important feature should result in greater model error than removing information about the value of a less important feature.
+Thus, we can assess a ranking by measuring the error of the model when the features the ranking deems most important are removed.
+More precisely, for a ranking $\pi$ and for each $k=1,2,...$ we measure the RMSE of $\xi(x,[d]\setminus\pi(x)[1..k])$ over all $x$ from the test dataset, where $\xi(\cdot,S)$ is a function giving model predictions based only on features from $S$.
+We have two such functions $\xi$ (and thus two different methods of comparing rankings) defined below.
+The first method is feature randomising.
+We define $\xi_{random}(x,S):=\mathbb{E}[m(X)]$ where $m$ is the model and $X$ is a random variable with $X_{i}\equiv x_{i}$ for $i\in S$ and $X_{j}$ has the distribution of the $j$-th feature over the test set for $j\not\in S$.
+In practice, the expected value is approximated by sampling $m(X)$ 100 times (where each feature is sampled independently) and taking the average.
+The other method follows the remove-and-retrain approach described in (Hooker et al.
+2019).
+For $k=1,2,3,...$ and for each subset of features $S\subset[d]$ such that $|S|=d-k$ we train a XGBoost model where we only use the features in $S$.
+That is, the $k$ features that are not in $S$ are not available for the training of the model.
+This way, we obtain a family of models $(m_{S})_{S\subset[d]}$ where each model $m_{S}$ gives predictions based only on values of features in $S$.
+We define $\xi_{retrain}(x,S):=m_{S}(x)$.
+The comparison was performed for $k=1,2,3$.
+The results are shown in Tables in the Appendix.
+The results for the retraining method suggest that $\pi_{\text{SHAP}}$ picks the most important features more accurately than $\pi_{\mathrm{PG}}^{\sigma}$.
+The results for the randomising method show no significant difference between $\pi_{\text{SHAP}}$ and $\pi_{\mathrm{PG}}^{\sigma}$ for Red Wine Quality and Parkinson Telemonitoring models.
+However, for the Housing model we see that $\pi_{\mathrm{PG}}^{\sigma}$ with $\sigma\in\{0.1,0.3\}$ give better results than $\pi_{\text{SHAP}}$.
+In general, this comparison doesn’t give a simple conclusion that one ranking method is clearly better than the other.
+Each of them can show advantage over the other depending on the dataset and assessment method.
+Conclusion
+In this work, we proposed a cleanly defined faithfulness metric $\mathrm{PGI}^{2}$ that can be computed exactly in polynomial time on tree ensemble models.
+We designed a quadratic algorithm for that.
+Our experimental evaluation showed that the algorithm is numerically stable and superior to Monte Carlo-based methods if very accurate results are desired.
+We also proposed a natural feature ranking method inspired by $\mathrm{PGI}^{2}$ optimization.
+The method generally identifies different most important features than the state-of-the-art SHAP explainer while offering similar performance wrt.
+different faithfulness metrics.
+The $\mathrm{PG}^{2}$-based ranking may thus offer a viable alternative for assessing feature importance in tree-based models.
+Acknowledgments
+Partially supported by the ERC PoC grant EXALT no 101082299 and the National Science Centre (NCN) grant no. 2020/37/B/ST6/04179.
 Appendix A Models training
-The grid search was performed by using class $GridSearchCV$ from library XGBoost [6] with numbers of folds in cross-validation set to $5$.
-Additionally to parameters described in 4.2, the grid search included also parameters with the following possible values:
+The grid search was performed by using class $GridSearchCV$ from library XGBoost (Chen and Guestrin 2016) to minimize squared error, with numbers of folds in cross-validation set to $5$.
+Additionally to parameters described in Models, the grid search included also parameters with the following possible values:
+  • $max\_depth$ with possible values in $\{1,2,3,4\}$
   • $eta$ with possible values in $\{0.01,0.1,0.2,0.3,0.4,0.5,0.6,0.7,0.8,0.9\}$
   • $subsample$ with possible values in ${0.01,0.1,0.2,0.3,0.4,0.5,0.6,0.7,0.8,0.9}$
-Appendix B Additional experiments results
-B.1 Results from precision experiments
+The chosen parameters for models were:
+  • Single tree model for Wine Quality dataset – $eta=0.9$, $subsample=0.9$and $max\_depth=4$
+  • Bigger model for Wine Quality dataset – $eta=0.2$, $subsample=0.9$, $max\_depth=4$ and $40$ trees
+  • Single tree model for Californian Housing dataset – $eta=0.9$, $subsample=0.9$ and $max\_depth=3$
+  • Bigger model for Californian Housing dataset – $eta=0.2$, $subsample=0.9$, $max\_depth=4$ $40$ trees
+  • Single tree model for Parkinson Telemonitoring dataset – $eta=0.9$ and $subsample=0.9$, $max\_depth=3$
+  • Bigger model for Parkinson Telemonitoring dataset – $eta=0.3$, $subsample=0.8$ and $max\_depth=4$
+Appendix B Quasi Monte Carlo
+In our implementation of the Quasi-Monte Carlo method, we utilized the quasi-random generator provided by the SciPy library (Virtanen et al.
+2020).
+Specifically, we employed the Halton sequence (Halton 1960) for generating our quasi-random points.
+This choice was motivated by the Halton sequence’s flexibility, as it does not necessitate a sample size that is a power of two $(2^{n})$, being a better fit for experimental designs.
+To transform the uniform distribution produced by the Halton sequence into the required normal distribution, we applied the inverse transform method.
+It is worth noting that while this transformation is effective, it does result in some loss of the low-discrepancy properties inherent to the original Halton sequence.
+Appendix C Additional experiments results
+Results from precision experiments
 Figure 2: Single tree model for Wine Quality dataset
@@
 Figure 5: Bigger model for Californian Housing dataset
-B.2 Results from PGI comparison experiment
-Figure 11: Single tree model for the Wine Quality dataset.
-Figure 12: Bigger model for the Wine Quality dataset.
-Figure 13: Single tree model for Californian Housing dataset
-Figure 14: Bigger model for Californian Housing dataset
+Figure 6: Single tree model for Parkinson Telemonitoring dataset
+Figure 7: Bigger model for Parkinson Telemonitoring dataset
+Results from feature removing experiment
+Table 4: Rankings comparison wrt. feature randomising metric
+$k$ SHAP $\sigma$ = $0.1$ $\sigma$ = $0.3$ $\sigma$ = $1.0$
+Bigger Red Wine Quality model
+1 0.944 0.922 0.923 0.919
+2 0.893 0.902 0.881 0.891
+3 0.858 0.89 0.87 0.868
+Bigger California Housing model
+1 1.32e+05 1.412e+05 1.379e+05 1.264e+05
+2 1.3e+05 1.354e+05 1.321e+05 1.263e+05
+3 1.282e+05 1.318e+05 1.287e+05 1.27e+05
+Bigger Parkinson Telemonitoring model
+1 8.87 8.96 8.99 8.945
+2 8.748 8.882 8.882 8.906
+3 8.585 8.777 8.809 8.784
+Table 5: Rankings comparison wrt. remove-and-retrain metric
+$k$ SHAP $\sigma$ = $0.1$ $\sigma$ = $0.3$ $\sigma$ = $1.0$
+Bigger PGI Red Wine Quality model
+1 0.67 0.621 0.637 0.667
+2 0.72 0.638 0.66 0.672
+3 0.722 0.652 0.665 0.711
+Bigger PGI California Housing model
+1 8.092e+04 7.187e+04 7.206e+04 7.485e+04
+2 8.913e+04 7.693e+04 8.062e+04 8.707e+04
+3 9.136e+04 8.267e+04 8.965e+04 9.665e+04
+Bigger PGI Parkinson Telemonitoring model
+1 7.216 6.745 6.863 7.02
+2 7.388 6.951 7.138 7.292
+3 7.55 7.021 7.282 7.45
+Results from PGI comparison experiment
+Table 6: Greedy $\mathrm{PG}^{2}$ vs SHAP rankings wrt. $\mathrm{PGI}^{2}$.
+$\sigma^{\prime}$ SHAP $\sigma$ = $0.1$ $\sigma$ = $0.3$ $\sigma$ = $1.0$
+Single Red Wine Quality model
+0.1 0.021 0.023 0.023 0.023
+0.3 0.077 0.095 0.096 0.096
+1.0 0.151 0.174 0.185 0.185
+Single California Housing model
+0.1 3.267e+08 3.987e+08 3.987e+08 3.987e+08
+0.3 1.421e+09 1.582e+09 1.597e+09 1.597e+09
+1.0 5.921e+09 6.276e+09 6.451e+09 6.459e+09
+Single Parkinson Telemonitoring model
+0.1 1.084 1.084 1.084 1.084
+0.3 3.438 3.573 3.823 3.823
+1.0 6.143 7.322 8.444 8.449